OpenStudy (anonymous):

What is the slope of the line that passes through the points (–6,1) and (4,–4)? A. -2 B. 2 C. -½ D. ½

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=(-4-1)/(4--6) m=(-5)/(10) m=-1/2